OpenStudy (anonymous):

Write y = 1/6x + 5 in standard form using integers. A. –x – 6y = 30 B. 6x – y = 30 C. –x + 6y = 30 D. –x + 6y = 5

OpenStudy (anonymous):

I think the answer is C?

OpenStudy (lgbasallote):

\[y = \frac 16 x + 5\] multiply all terms by 6 \[\implies 6y = x + 30\] subtract x from both sides \[\boxed{\implies -x + 6y = 30}\] so..answer's C

OpenStudy (anonymous):

Thank you!!

OpenStudy (lgbasallote):

welcome